OpenStudy (anonymous):

Aisha just received a $40 paycheck from her new job. She spends some of it buying music online and saves the rest in a bank account. Her savings is given by f(x) = 40 ­ 1.25x, where x is the number of songs she downloads at $1.25 per song. a. Graph the function.

OpenStudy (anonymous):

edit: f(x) = 40 - 1.25x

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (usukidoll):

put it in y = mx+b form f(x) also means y we just need rearranging for the formula

OpenStudy (anonymous):

y = -1.25x + 40

OpenStudy (usukidoll):

yup... now we just graph that... hmm..... not by hand .... with desmos online grapher

OpenStudy (anonymous):

i have to take a pic of it to send it through online, so i gotta do it by hand lol

OpenStudy (usukidoll):

if we chose x = 0 then y = 40 and that is crazy to graph manually

OpenStudy (usukidoll):

oh wow XD good luck XD

OpenStudy (anonymous):

lol, can u tell me what goes on the x value line and the y value so i can graph?

OpenStudy (usukidoll):

I think it's torture .... when x = 0 y = 40 so by graph paper you will have to combine sheets

OpenStudy (usukidoll):

well we can just let x = 0,1,2,3... and then solve for y same process when x = -1,-2,-3 and then solve for y

OpenStudy (anonymous):

i mean, when i graph this manually, the y values will be 0 to 40 and the x values will be..1 2 3 etc?

OpenStudy (usukidoll):

we can choose our x ... that's why x can be 0,1,2,3. we plug those values into the equation so we can solve for y when x = 0 y = 1.25(0) +40 y=0+40 = 40 so (0,40) is a point on the graph

OpenStudy (usukidoll):

when x = 1 y=1.25(1) +40 = 1.25+ 40 = 41.25 so (1,41.25)

OpenStudy (usukidoll):

that's torture to draw manually

OpenStudy (anonymous):

hmm yea >.< lol i'll see what i can do, thanks!

OpenStudy (usukidoll):

ugh .. I forgot the negative ok that didn't make an impact when x = 0 but it does when x = 1 so y=-1.25+40 = 38.75 (1, 38.75) that makes better sense.

OpenStudy (usukidoll):

damn I was half drowsy when I was doing the x = 1 xD

OpenStudy (anonymous):

 this is what im kinda confused about

OpenStudy (anonymous):

i plot 40 on the y axis, and i gotta go down -1.25 cuz thats the slope, but what numbers do i put on the x axis

OpenStudy (usukidoll):

if your graphing by 5's... it should be uniformed..